OpenStudy (anonymous):

Larry uses his credit card to purchase a new video game system for $445.07. He can pay off up to $200 per month. The card has an annual rate of 13.6% compounded monthly. How much total interest will he pay? $8.47 $10.21 $53.54 $3.32

OpenStudy (kropot72):

Assume Larry pays off $200 each month. The interest at the end of month 1 is \[445.07\times \frac{0.136}{12}=$\ can\ you\ calculate?\]

OpenStudy (kropot72):

@Fenetre135 Are you there?

OpenStudy (anonymous):

yeah just give me a sec :)

OpenStudy (anonymous):

5.04?

OpenStudy (kropot72):

Good work! The balance at the end of the first month after adding interest and subtracting the payment is \[445.07+5.04-200=you\ can\ calculate\]

OpenStudy (anonymous):

250.11

OpenStudy (kropot72):

Correct again! The interest at the end of the second month is \[250.11\times \frac{0.136}{12}=$\ ?\]

OpenStudy (anonymous):

2.83

OpenStudy (kropot72):

Good! The balance at the end of the second month after adding interest and subtracting the payment is \[250.11+2.83-200=$\ ?\]

OpenStudy (anonymous):

52.94

OpenStudy (kropot72):

Good! We are heading to one of the choices. The interest at the end of the third month is \[52.94\times \frac{0.136}{12}=$\ x\] When you have found the value of x do the following addition to find the total interest \[$5.04+$2.83+$\ x=$\ you\ can\ calculate\]

OpenStudy (anonymous):

8.469

OpenStudy (anonymous):

thank you for your help!

OpenStudy (kropot72):

So the correct choice is $8.47. You're welcome :)